Find the general solution of the differential eq. y^2/2+2ye^t+(y+e^t)dy/dt=0
1
Expert's answer
2019-09-27T09:35:34-0400

let's rewrite the equation in the form

"(\frac{y^2}{2}+2ye^t)dt+(y+e^t)dy=0"

This is not an exact equation, because

"\frac{\partial}{\partial y}(\frac{y^2}{2}+2ye^t)=y+2e^t\ne e^t=\frac{\partial}{\partial t}(y+e^t)"

Therefore we need to find an integrating factor "\mu" , such that

"\frac{\partial}{\partial y}\mu(t)(\frac{y^2}{2}+2ye^t)=\frac{\partial}{\partial t}\mu(t)(y+e^t)"

"\mu(t)(y+2e^t)=\frac{d\mu}{dt}(y+e^t)+\mu(t)e^t"

"\mu(t)(y+e^t)=\frac{d\mu}{dt}(y+e^t)"

"\mu(t)=\frac{d\mu}{dt}"

"\mu=e^t"

Then

"(e^t\frac{y^2}{2}+2ye^{2t})dt+(e^ty+e^{2t})dy=0"

is an exact equation

Let's integrate

"\int(e^t\frac{y^2}{2}+2ye^{2t})dt=e^t\frac{y^2}{2}+4ye^{2t}+C(y)"

Then

"\frac{d}{dy}\left(e^t\frac{y^2}{2}+4ye^{2t}+C(y)\right)\\=ye^t+4e^{2t}+C'=ye^t+e^{2t}"

"C'=-3e^{2t}"

"C=-\frac{3e^{2t}}{2}"

Therefore, the solution of the equation will be

"e^t\frac{y^2}{2}+4ye^{2t}-\frac{3e^{2t}}{2}=const"